a. In a​ poll,

44​%

of

1381

surveyed adults said that

pesticides arepesticides are

"very harmful""very harmful".

What is the actual number of adults who said that

pesticides arepesticides are

"very harmful" question mark nbsp"very harmful"? b. Among the

1381

surveyed​ adults,

1364

said that

pesticides arepesticides are

"not at all harmful""not at all harmful".

What is the percentage of people who chose "not at all harmful" question mark nbsp"not at all harmful"?

a. What is the actual number of adults who said that

pesticides arepesticides are

"very harmful" question mark nbsp"very harmful"?

nothing

​(Round to the nearest integer as​ needed.)b. What is the percentage of people who chose

​"not at all harmfulnot at all harmful​"?

nothing​%

​(Round to the nearest tenth as​ needed.)
Answer #1

Answer

(A) given proportion is p = 0.44 and sample size is n = 1381

Number of people who said “very harmful” = p*n

setting the n and p values, we get

= 0.44*1381

= 607.64

= 608(rounded to nearest)

(B) total number of people n = 1381

number of people who said not at all harmful (x)= 1364

required percentage = (x/n)*100

setting the values, we get

=(1363/1381)*100

= 98.8%